OpenStudy (anonymous):

A right triangle’s hypotenuse is 20 cm long. What is the length of the side opposite a 60° angle? Give your answer to the nearest tenth of a centimeter. 40 18.5 20 17.3

OpenStudy (anonymous):

given this is a 30-60-90 triangle, the short side (opposite the 30) is always half the hyp, so 20/2 = 10 cm. Once you get this, simply apply pythagorus' formula : sqrt(20^2 - 10^2) = b sqrt300 = b And I don't have a calculator for this last step

OpenStudy (anonymous):

ohk thanks

jhonyy9 (jhonyy9):

of you can ising sin 60 = opposite side / hypotenuse - you know sin 60 =(sqrt3)/2 - so hence from these you can calcule the lenght of opposite side - so what will be : lenght of opposite side = ((sqrt3)/2) * 20 = 1,73 *10 = 17,3